# Randomized Masked Finetuning: An Efficient Way to Mitigate Memorization of PIIs in LLMs

### v1

@@ Abstract
The current literature on memorization in Natural Language Models, especially Large Language Models (LLMs), poses severe security and privacy risks, as models tend to memorize personally identifying information (PIIs) from training data. We introduce Randomized Masked Fine-Tuning (RMFT), a novel privacy-preserving fine-tuning technique that reduces PII memorization while minimizing performance impact. Using the Enron Email Dataset, we demonstrate that RMFT achieves an 80.81% reduction in Total Extraction Rate and 80.17% reduction in Seen Extraction Rate compared to baseline fine-tuning, outperforming deduplication methods while maintaining only a 5.73% increase in perplexity. We present MaxTER, a Pareto-optimal evaluation framework for assessing privacy-utility tradeoffs, and show the performance of RMFT vs Deduplication by Area Under The Response Curve (AURC) metric.
@@ 1 Introduction
The current literature on memorization in Natural Language Models, especially Large Language Models (LLMs), poses a severe security and privacy risk  (carlini2021extractingtrainingdatalarge; yang2024memorization), as models tend to memorize pieces of personally identifying information (PIIs) that occur in pre-training or fine-tuning datasets and can later regurgitate the exact PIIs when prompted. This effect arises from models’ ability to identify token sequences in PIIs that occur frequently, which leads to memorization of the exact PII as compared to generalization and generation when prompted. This was explored by Kandpal and team (kandpal-etal-2024-user) in their study on deduplication training and how it mitigates privacy risks. They and others  (10.1162/TACL.a.49) have observed that token sequences that were duplicated multiple times in the dataset also tend to be memorized more often than token sequences that appeared a smaller number of times or only once.
Our preliminary examination on a GPT2-XL tokenizer revealed that the tokenizer considered a PII as a composition of several tokens. This means that PIIs that occurred multiple times tend to be memorized as a token sequence by the LLM. Therefore, we move forward with the assumption that the N-Gram appearances of PIIs directly impact the ability of an LLM to memorize the PII.
Keeping this in mind, we devised a novel training technique that reduces the number of times a PII appears in the dataset while keeping the original PII intact for future downstream tasks. This technique is called Randomized Masked Fine-Tuning (RMFT). As the name suggests, we mask the PII appearances with a randomized token sequence generation that closely mimics the PII being masked. Using this method, we were able to reduce the number of appearances of certain PIIs from thousands down to a single appearance.
We fine-tune the model with this modified dataset and then test the updated model on the same prompts. We show that RMFT is highly efficient in reducing memorization in LLMs compared to baseline fine-tuning with no dataset modifications, with a minor impact on model performance as measured using perplexity on the WikiText2 benchmark dataset.
We then compare our technique with the deduplication technique proposed by Katherine Lee and team. (lee2022deduplicating) and create a comprehensive suite of evaluations on PII extraction, model performance, and training efficiency. Using our results, we show how RMFT outperforms deduplicated models when it comes to PII extraction on the same prompts.
@@ 1.1 Contributions
  • Comprehensive evaluation metrics: Including Total Extraction Rate (TER) and Seen Extraction Rate (SER) for measuring memorization, and Mean Delta Perplexity (MDP) for performance assessment.
  • MaxTER: A new optimization function using Pareto frontier analysis to evaluate the memorization-versus-performance tradeoff.
  • Empirical validation: Demonstrating RMFT’s superiority over deduplication on the Enron Email Dataset with GPT2-XL.
@@ 2.1 PII Memorization and Leakage
Prior work on Personally Identifiable Information (PII) memorization and leakage has revealed a critical, inherent privacy vulnerability in large language models (LLMs). Multiple earlier studies successfully retrieved training data, including PII, from publicly available GPT-2 models (carlini2021extractingtrainingdatalarge; nasr2023scalableextractiontrainingdata; schwarzschild2024rethinkingllmmemorizationlens). Huang and team (huang-etal-2022-large) formally coined the term memorization in 2022 during their work on extracting personal information from pre-trained LLMs. Borkar’s study (borkar2023can) worked extensively on data leakage related to law-based datasets and accredited unlearning as a possible solution to memorization. Extending this concept, Borkar ad team  (borkar2025privacyrippleeffectsadding) introduced the detailed concepts of assisted memorization, retained memorization, forgotten memorization, and immediate memorization. Furthermore, multiple studies have focused on the memorization of token sequences, phrases, and passages and the capacity of LLMs to reproduce them exactly as available in the training data (satvaty2025undesirablememorizationlargelanguage; hartmann2023sokmemorizationgeneralpurposelarge; wang2025generalizationvsmemorizationtracing).
@@ 2.2 Mitigation Techniques
There are multiple mitigation techniques that aim at reducing the memorization of PIIs and token sequences in LLMs while maintaining minimal impact on model performance. Unlearning (borkar2025privacyrippleeffectsadding; chen2023unlearnwantforgetefficient; si2023knowledgeunlearningllmstasks; tamirisa2024toward) and deduplication (kandpal-etal-2024-user; lee2022deduplicating) are often cited as the most efficient mitigation techniques currently. However, a study by To and Le  (to2025harrypotterhereprobing) indicated that unlearning can be inefficient and introduced a framework called LURK, which evaluates unlearned models for retained memorization of unlearned data. Kandpal  (pmlr-v162-kandpal22a) accredited differential privacy as an efficient mitigation technique for membership inference. In a comparative study, Sakarvadia and team  (sakarvadia2025mitigatingmemorizationlanguagemodels) compared 11 different techniques used to combat memorization and introduced balanced subnet, a model modification-based method. Separately, Suri and team  (suri2025mitigatingmemorizationllmsusing) proposed activation steering on privacy neurons as an efficient mitigation method.
2.3 Privacy vs Utility in LLMs
The consistent need to balance model performance with privacy guarantees has led to research in the quantification of privacy and utility in LLMs. Sun and team (10.1145/3696410.3714531) focused on researching the effectiveness of privacy-preserving algorithms in LLMs, concluding that the CusText+ technique was optimal for robust privacy preservation. Das and team (das2025revisitingprivacyutilityefficiency) attempted to redefine privacy and utility specifically in the context of fine-tuning LLMs, evaluating different fine-tuning techniques such as differential privacy and LoRA. Pasch and team (pasch-cha-2025-balancing) studied anonymizations and utility in LLMs, concluding that anonymizations can reduce privacy leakage by 97-99% while only causing a 10% degradation in the model. These results serve as direct motivation for our work, as we choose to mask PIIs anonymously to reduce PII leakage while minimizing the impact on LLM performance.
Although existing studies have explored the impact of privacy-preserving techniques on LLM utility, they typically evaluate privacy and utility considerations separately rather than within an integrated analytical framework. To the best of our knowledge, this work represents the first systematic investigation of the privacy-utility tradeoff for LLM mitigation techniques under a unified framework that jointly characterizes both dimensions.
@@ 3 Dataset Used
We utilize a raw version of the Enron Email Dataset (shetty2004enron) to construct our training, validation, and testing splits. The dataset is divided into three partitions following a 100:10:1 ratio—the training set contains 10,000 datapoints, the validation set 1,000 datapoints, and the testing set 100 datapoints, all randomly sampled. Each datapoint corresponds to an individual email sample containing one or more personally identifiable email addresses.
@@
  • Memorization Evaluation via Generation: We employ a set of 16,000 English prompts sampled from a publicly available web crawl. These prompts are designed to mimic realistic, open-ended user queries rather than explicitly targeted extraction prompts, ensuring the evaluation reflects general LLM usage patterns.
@@
For randomized masking, we construct a small, curated list of commonly used domain names and top-level domains (TLDs), publicly available on our GitHub repository  (joshi2024rmft_code). Each list contains approximately 15 examples, used to generate realistic but anonymized email variants during masking.
@@ 3.1 Exploratory Data Analysis
Before applying any dataset modifications, we conducted an exploratory data analysis (EDA) on the Enron Email dataset to better understand the statistical distribution of personally identifiable email addresses. Specifically, we aimed to quantify how often certain PIIs recur across the dataset, since our hypothesis—consistent with Kandpal and team  (kandpal-etal-2024-user) —is that frequent token repetition correlates with higher memorization likelihood in LLMs.
Figure 1 illustrates the distribution of the top ten most frequent email addresses across both the training and testing subsets. Some addresses appear nearly 900 times among the 10,000 datapoints in the training split, indicating strong duplication patterns. Such high-frequency occurrences are particularly problematic for privacy, as repeated exposure during fine-tuning can lead models to memorize entire token sequences corresponding to these PIIs. Additionally, we observed that several email addresses appear in both training and testing subsets, highlighting potential data leakage across splits.
These findings directly motivated our two dataset-level interventions: (1) Randomized Masked Fine-Tuning (RMFT), which reduces recurrence through randomized replacements while preserving structure; and (2) Deduplication, which explicitly removes repeated headers to prevent repeated exposure to the same PII tokens.
(a) Training set
(b) Testing set
Figure 1: Distribution of the top ten most frequent email addresses in training and testing sets.
@@ 4 Methodology
We perform experiments to test the effectiveness of randomized masking in preventing PII memorization. As we have taken our inspiration from Kandpal and team’s results on N-Gram deduplication, we also compare our results with an adapted version of deduplication for PIIs. Figure 2 depicts a summarized version of methodologies used in randomized masked fine-tuning and deduplication.
Figure 2: Methodology used for Randomized Masked Finetuning vs Deduplication. This diagram depicts how the index tables are formulated for either technique and how the updated index tables differ. If we have $N$ datapoints, $k$ emails occurring on each datapoint on average, and $u$ unique emails, then the number of times each email appears in the original index table is $Nk/u$. In RMFT, the updated index table maintains NkNk rows, but each email now appears only once and the number of unique emails increases. In deduplication, the updated index table has just $u$ rows with each email appearing once.
@@
Over both phases we are fine-tuning our model for 3 epochs, checkpointing the LLM at every 10% training steps, utilizing a fixed learning rate of 0.00002 over all three training methods, as inspired by Borkar’s work  (borkar2025privacyrippleeffectsadding). We also save the data seen until every checkpoint and compare checkpoint-based metric calculation for the different fine-tuning techniques. For generation experiments, greedy decoding with no sampling is used, setting the temperature to 0, top_p to 1.0, top_k to 0, and deterministic beam search switched off. The average TER, average SER, and average mean delta perplexity are calculated as the average respective values over the 30 checkpoints.
@@ 4.1.2 Deduplication Training
We adapt deduplication for PII-aware context, closely following the structure defined by Lee and team  (lee2022deduplicating). Lee and team define deduplication as dropping datapoints that contain token sequences of 400 tokens or more that are repeated multiple times. We adapt it to PII-aware context, where we drop datapoints that have a repeating email address.
@@ 4.2 Metrics Used
We use two metrics to calculate and compare memorization: Total Extraction Rate (TER) and Seen Extraction Rate (SER). We utilize Mean Delta Perplexity (MDP) to compare the performance of the training techniques.
@@ 4.2.1 Memorization Metrics
Total Extraction Rate is defined as the number of email addresses leaked at a checkpoint divided by the total number of email addresses available in the original set. Leaked email addresses are extracted from generated text using regex pattern matching for the format [first_name].[last_name]@[domain].[tld], with matches validated against the original email set.
TER is calculated as:
@@ 5.3 Time Complexity Analysis
We assume $N$ datapoints for training, each datapoint has on average $k$ PIIs appearing, the length of each datapoint is $l$, and the number of unique emails is $u$. The total number of PIIs occurring in the whole dataset would be $N\cdot k$.
@@ Appendix B MaxTER Evaluation on Different Datasets
We present evaluations of RMFT versus deduplication via the MaxTER function on three datasets: Enron held-out test dataset, WikiText2 test dataset  (radford2019language), and randomly collected web crawl prompt dataset. Table 2 summarizes all results.
@@
Appendix C Time Complexity Analysis
For this analysis, we assume that we have $N$ datapoints, each datapoint being $l$ length long, each datapoint having $k$ emails on average, and there are $u$ unique emails all over the dataset.
C.1 Randomized Masked Finetuning
We need to sort all email addresses fetched in the index table based on their occurrence. Each email appears $\frac{N\cdot k}{u}$ times. Therefore, using quick sort the time complexity for sorting becomes:
$TC_{sorting}$ $=\frac{N\cdot k}{u}\log\left(\frac{N\cdot k}{u}\right)$ (7)
$=\frac{N\cdot k}{u}(\log N+\log k-\log u)$ (8)
As we repeat this for $u$ unique emails:
$TC_{sorting}=u\cdot\frac{N\cdot k}{u}(\log N+\log k-\log u)$ (9)
In our case the number of unique emails is of higher magnitude than number of datapoints $N$ and average email appearance $k$. Therefore:
$TC_{sorting}=O(N\cdot k\cdot|\log u|)$ (10)
The randomization and replacement of PIIs occur in $O(1)$ for $N\cdot k$ emails:
$TC_{RR}=O(N\cdot k)$ (11)
The time complexity of RMFT is:
$\mathbf{TC_{RMFT}=O(N\cdot k\cdot|\log u|)}$ (12)
C.2 Deduplication
The deduplicator takes $O(1)$ time for deduplication and we carry it out for $u$ unique emails:
$TC_{Deduplicator}=O(u)$ (13)
We have $N$ headers on average and sorting them gives:
$TC_{HeaderSorting}=O(N\log N)$ (14)
The remerging takes $O(N)$. Therefore:
$TC_{Deduplication}=O(u)+O(N\log N)+O(N)$ (15)
We report the dominating term:
$\mathbf{TC_{Deduplication}=O(N\log N)}$ (16)

### v2

@@ Abstract
Language models are known to memorize training data, which they can later regurgitate when prompted in specific ways. Memorization may result in severe privacy concerns, especially when training datasets contain sensitive data like personally identifiable information (PII). We introduce Randomized Masked Fine-Tuning (RMFT), a novel privacy-preserving fine-tuning technique that reduces PII memorization while minimizing performance impact. Using the Enron Email Dataset, we demonstrate that RMFT achieves an 80.81% reduction in Total Extraction Rate and 80.17% reduction in Seen Extraction Rate compared to baseline fine-tuning, outperforming deduplication methods while maintaining only a 5.73% increase in perplexity. We present MaxTER, a Pareto-optimal evaluation framework for assessing privacy-utility tradeoffs, and show the performance of RMFT vs Deduplication by Area Under The Response Curve (AURC) metric.
@@ 1 Introduction
Language models are known to memorize training data, which they can later regurgitate when prompted in specific ways (Carlini et al., 2021). Memorization may result in severe privacy concerns, especially when training datasets contain sensitive data like personally identifiable information (PII) (Yang et al., 2024). Prior studies show that memorization in language models is largely driven by data duplication (Lee et al., 2022; Kandpal et al., 2022)—token sequences that appear multiple times in the dataset tend to be memorized more often than sequences appearing once or infrequently (Kandpal et al., 2024; Slack and Al Moubayed, 2025). Therefore, we move forward with the assumption that a higher number of duplicates for a specific PII increases the risk of memorization.
Keeping this in mind, we devised a novel fine-tuning technique that reduces the number of times a particular PII appears in the dataset while keeping the original PII intact for future downstream tasks. This technique is called Randomized Masked Fine-Tuning (RMFT). In our study, we focus on email addresses as the PII of interest. RMFT identifies all occurrences of each email in the dataset and preserves only the first occurrence in its original form. All subsequent duplicate occurrences are masked by generating structurally similar email addresses using randomly selected components—firstnames, lastnames, and domains—that closely mimic the original email’s format. Using this method, we reduce individual email occurrences from hundreds or thousands down to a single appearance.
We fine-tune the model with this modified dataset and evaluate memorization using an extractable memorization setup, where we prompt the model with the same set of randomly sampled prompts used for the baseline fine-tuned model to ensure uniformity across experiments. Our results demonstrate that RMFT is highly effective in reducing memorization: we achieve an 80.81% reduction in Total Extraction Rate (TER) and an 80.17% reduction in Seen Extraction Rate (SER) compared to baseline fine-tuning with no dataset modifications. This substantial privacy improvement comes with minimal impact on model performance—RMFT incurs only a 5.73% increase in perplexity as measured on the Enron Testing Split.Validation on GPT-Neo-1.3B demonstrates similar results (80.42% TER reduction, 7.08% perplexity increase), confirming generalizability across architectures (Appendix C).
To benchmark RMFT’s effectiveness, we adapt the deduplication methodology from Lee et al. (2022) to our PII-specific context by removing datapoints containing duplicate email addresses rather than duplicate text sequences. We then conduct a comprehensive evaluation comparing RMFT and deduplication across three dimensions: email extraction rates, model performance, and training efficiency. Our results demonstrate that RMFT achieves better privacy-utility tradeoff as compared to deduplication. However, deduplication is computationally more efficient, requiring less preprocessing time than RMFT’s masking procedure.
@@ 1.1 Contributions
  • Novel Privacy-Utility Evaluation Framework: We introduce MaxTER, a Pareto-optimal analysis framework that jointly characterizes the privacy-utility tradeoff in memorization mitigation techniques. Adapting extraction methodologies from Carlini et al. (2021) and Nasr et al. (2023), we define Total Extraction Rate (TER) and Seen Extraction Rate (SER) as privacy metrics (Section 4.2.1), and Mean Delta Perplexity (MDP) as our utility metric (Section 4.2.2). For a given perplexity threshold $\tau$, MaxTER identifies the checkpoint that achieves maximum TER reduction, enabling practitioners to select optimal privacy-utility operating points. We quantify overall framework efficiency using Area Under the Response Curve (AURC) to systematically compare RMFT and deduplication across privacy-utility tradeoff spaces.
@@ 2.1 PII Memorization and Leakage
Large language models have been shown to memorize and leak personally identifiable information (PII) from training data, with successful extraction demonstrated on publicly available models like GPT-2 (Carlini et al., 2021; Nasr et al., 2023; Huang et al., 2022). Beyond individual PII extraction, research has characterized how memorization manifests throughout training, identifying patterns such as assisted, retained, forgotten, and immediate memorization (Borkar et al., 2025). Studies have also examined how models reproduce token sequences, phrases, and entire passages verbatim (Hartmann et al., 2023; Wang et al., 2025).
@@ 2.2 Mitigation Techniques
Current approaches to reducing memorization include unlearning—removing specific data post-training (Chen and Yang, 2023; Si et al., 2023)—and deduplication, which removes repeated sequences before training (Kandpal et al., 2024; Lee et al., 2022). However, unlearning can be inefficient and may not fully eliminate memorization (To and Le, 2025). Alternative techniques include differential privacy for membership inference protection (Kandpal et al., 2022), model architecture modifications (Sakarvadia et al., 2025), and activation steering on privacy-sensitive neurons (Suri et al., 2025).
2.3 Deduplication as a Privacy-Preserving Technique
Among mitigation techniques, deduplication has emerged as particularly effective and practical for reducing memorization. Lee et al. (2022) demonstrated that exact deduplication—removing training sequences appearing verbatim multiple times—substantially reduces memorization rates while improving model performance on downstream tasks. Extensions beyond exact matching, such as semantic deduplication (Abbas et al., 2023), demonstrate that removing semantically similar sequences can further reduce memorization while preserving dataset diversity. Tirumala et al. (2022) confirmed that deduplication consistently reduces memorization throughout training dynamics without sacrificing model capability, while Kandpal et al. (2022) established the underlying causal mechanism: token sequences appearing more frequently in training data exhibit proportionally higher extraction rates. These findings collectively establish deduplication as a principled and effective approach to mitigating memorization.
However, standard deduplication operates at the document or sequence level—typically removing passages of 50+ tokens that appear multiple times—and does not specifically target structured PII such as email addresses, which may be embedded within otherwise unique contexts. This gap motivates our work: we adapt deduplication to PII-aware scenarios by removing datapoints containing duplicate email addresses, and we introduce RMFT as an alternative that reduces duplication while preserving dataset size and linguistic diversity. Given deduplication’s established effectiveness and its conceptual similarity to our approach, we use the adapted version as explained in 4.1.2 as our primary baseline for comparison.
2.4 Privacy-Utility Tradeoffs
Balancing privacy preservation with model utility remains a central challenge in LLM deployment. Recent work has evaluated various privacy-preserving techniques (Sun et al., 2025; Das et al., 2025), with promising results showing that anonymization can reduce privacy leakage by 97-99% while incurring only 10% performance degradation (Pasch and Cha, 2025). However, existing studies typically evaluate privacy and utility in isolation. Our work addresses this gap by proposing a unified framework that jointly characterizes privacy-utility tradeoffs for PII mitigation techniques.
@@ 3 Dataset Used
We utilize a raw version of the Enron Email Dataset (Shetty and Adibi, 2004) to construct our training, validation, and testing splits. The dataset is divided into three partitions following a 100:10:1 ratio—the training set contains 10,000 datapoints, the validation set 1,000 datapoints, and the testing set 100 datapoints, all randomly sampled. Each datapoint corresponds to an individual email sample containing one or more personally identifiable email addresses.
@@
  • Memorization Evaluation via Generation: We employ a set of 16,000 English prompts sampled from Common Crawl (CC-MAIN-2024-10, February 2024 snapshot). Prompts were extracted from the WARC file using a fixed random seed to ensure reproducibility and represent realistic, open-ended user queries rather than explicitly targeted extraction prompts. For generation, we use the following configuration: prompts are truncated to 50 tokens as input, greedy decoding, and a maximum generation length of 256 tokens (including the prompt). This setup ensures deterministic, reproducible extraction experiments across all checkpoints and training techniques. Email addresses are extracted from generated text using regex pattern matching, with matches validated against the original Enron dataset. Example prompts are provided in Appendix LABEL:app:prompts.
@@
For randomized masking, we construct a small, curated list of commonly used domain names and top-level domains (TLDs), publicly available on our GitHub repository (Anonymous, 2026). Each list contains approximately 15 examples, used to generate realistic but anonymized email variants during masking.
@@ 3.1 Exploratory Data Analysis
Before applying any dataset modifications, we conducted an exploratory data analysis (EDA) on the Enron Email dataset to better understand the statistical distribution of personally identifiable email addresses. Specifically, we aimed to quantify how often certain PIIs recur across the dataset, since our hypothesis—consistent with Kandpal et al. (2024)—is that frequent token repetition correlates with higher memorization likelihood in LLMs.
Figure 1 shows the frequency distribution on a log-log scale, revealing power-law (heavy-tailed) behavior in both training and testing sets. In the training set, the most frequent email appears 879 times among 10,000 datapoints, with frequencies declining smoothly across 24,480 unique email addresses. Such high-frequency occurrences are particularly problematic for privacy, as repeated exposure during fine-tuning can lead models to memorize entire token sequences corresponding to these email addresses.
This severe duplication directly motivated our two dataset-level interventions: (1) Randomized Masked Fine-Tuning (RMFT), which reduces recurrence through randomized replacements while preserving structure; and (2) Deduplication, which explicitly removes repeated headers to prevent repeated exposure to the same email tokens.
Figure 1: Email address frequency distributions on a log-log scale. Both training and testing sets exhibit power-law (heavy-tailed) distributions, with a small number of emails appearing hundreds of times while most appear once or twice. The most frequent email in the training set appears 879 times (rank 1), while the most frequent in the testing set appears 11 times. This severe duplication in training data motivates RMFT’s approach to reduce memorization risk by masking duplicate occurrences.
@@ 4 Methodology
We perform experiments to test the effectiveness of randomized masking in preventing PII memorization. As we have taken our inspiration from Kandpal results on N-Gram deduplication, we also compare our results with an adapted version of deduplication for PIIs. LABEL:fig:methodology depicts a summarized version of methodologies used in randomized masked fine-tuning and deduplication.
@@
Over both phases we are fine-tuning our model for 3 epochs, checkpointing the LLM at every 10% training steps, utilizing a fixed learning rate of 0.00002 over all three training methods, as in Borkar et al. (2025). We also save the data seen until every checkpoint and compare checkpoint-based metrics for the different fine-tuning techniques. For generation experiments, greedy decoding with no sampling is used, setting the temperature to 0, top_p to 1.0, top_k to 0, and deterministic beam search switched off. The average TER, average SER, and average mean delta perplexity are calculated as the average respective values over the 30 checkpoints.
@@ 4.1.2 Deduplication Training
We adapt deduplication for PII-aware context, closely following the structure defined by Lee et al. (2022). They define deduplication as dropping datapoints that contain token sequences of 400 tokens or more that are repeated multiple times. We adapt it to PII-aware context, where we drop datapoints that have a repeating email address.
@@ 4.2 Metrics Used
We adapt extraction rate methodologies from Carlini et al. (2021) and Nasr et al. (2023) to the PII context, defining two complementary metrics that measure memorization at different granularities: Total Extraction Rate (TER) for global PII extraction across the entire dataset, and Seen Extraction Rate (SER) for checkpoint-level PII extraction relative to training exposure. Borkar et al. (2025) similarly measure email address extraction rates over training sequences, which informed our metric design. For performance assessment, we utilize Mean Delta Perplexity (MDP) to compare training techniques.
@@ 4.2.1 Memorization Metrics
Total Extraction Rate is defined as the number of email addresses leaked at a checkpoint divided by the total number of email addresses available in the original set. Leaked email addresses are extracted from generated text using regex pattern matching for the format [first_name].[last_name]@[domain].[tld], with matches validated against the original email set. TER is calculated as:
@@ 5.3 Time Complexity Analysis
We assume $N$ datapoints for training, each datapoint has on average $k$ PIIs appearing, the length of each datapoint is $l$, and the number of unique emails is $u$. The total number of PIIs occurring in the whole dataset would be $N\cdot k$. For our experiments, N=10, [^0] N=10{,}000, $k\approx 6$, and u=24,480u=24{,}480.
@@ 8 Limitations and Future Work
We note that on out-of-distribution evaluation (WikiText2, Web Crawl), deduplication achieves higher AURC than RMFT (Appendix B). This occurs because RMFT is optimized for in-distribution privacy-preserving fine-tuning scenarios where deployment conditions match training distributions. Deduplication’s broader text-level approach may generalize better to unrelated domains, while RMFT’s PII-specific design excels at protecting sensitive information in the target domain.
@@
Our main evaluation focuses on GPT2-XL, with additional evaluation on GPT-Neo 1.3B (Appendix  C) demonstrate consistent and comparable performance. While the methodology is architecture-agnostic, validation across diverse model families (e.g., Llama, Qwen, Gemma, OLMo) would strengthen generalizability and is the focus of ongoing work.
@@ Appendix B MaxTER Evaluation on Different Datasets
We present evaluations of RMFT versus deduplication via the MaxTER function on three datasets: Enron held-out test dataset, WikiText2 test dataset  (Radford et al., 2019), and randomly collected web crawl prompt dataset. Table 2 summarizes all results.
@@
Appendix C Generalization to GPT-Neo-1.3B
To validate RMFT’s generalizability across model architectures, we repeated our experiments on GPT-Neo-1.3B, which differs from GPT-2-XL in both size (1.3B vs 1.5B parameters) and architecture details. We use identical experimental setup, training hyperparameters, and evaluation methodology as described in Section 4.
C.1 Results and Observations
Table 3 compares RMFT performance across GPT-2-XL and GPT-Neo-1.3B. Metrics are averaged over 30 checkpoints unless otherwise stated. Extraction rates are percentages.
Table 3: Comparison of RMFT performance on GPT-2-XL vs GPT-Neo-1.3B. Both models show consistent memorization reduction with minimal performance impact.
Model Avg TER Avg SER Avg PPL PPL
Reduction Reduction Increase
GPT-2-XL (1.5B) 80.81% 80.17% 5.73% 6.798
GPT-Neo (1.3B) 80.42% 80.21% 7.08% 5.156
C.1.1 Extraction Rates
RMFT significantly reduces memorization relative to baseline fine-tuning on GPT-Neo. RMFT achieves an 80.42% reduction in Total Extraction Rate (TER) and an 80.21% reduction in Seen Extraction Rate (SER) compared to the baseline. These results are highly consistent with our GPT-2-XL findings (80.81% TER reduction and 80.17% SER reduction), demonstrating that RMFT’s effectiveness generalizes across model architectures.
At the final checkpoint, RMFT achieves an 82.8% reduction in TER and 82.87% reduction in SER. These final checkpoint results are comparable to GPT-2-XL’s 87.62% TER reduction and 86.0% SER reduction, showing consistent privacy protection across both models.
Figure 7 compares the extraction rates across training checkpoints for both GPT-2-XL and GPT-Neo-1.3B.
(a) TER for GPT-2-XL
(b) TER for GPT-Neo-1.3B
(c) SER for GPT-2-XL
(d) SER for GPT-Neo-1.3B
Figure 7: Comparison of extraction rates across training checkpoints for GPT-2-XL and GPT-Neo-1.3B. (a) Total Extraction Rate (TER) for GPT-2-XL, (b) TER for GPT-Neo-1.3B, (c) Seen Extraction Rate (SER) for GPT-2-XL, and (d) SER for GPT-Neo-1.3B. Both models exhibit similar trends, with RMFT consistently achieving lower extraction rates than baseline across all checkpoints. Lower values indicate fewer emails memorized.
C.1.2 Mean Delta Perplexity
RMFT achieves an average perplexity increase of 7.08% relative to the baseline on GPT-Neo. While this is slightly higher than GPT-2-XL’s 5.73% increase, it remains within acceptable bounds for privacy-preserving fine-tuning. The consistency in memorization reduction (approx. 80%) across both models, despite minor variations in perplexity impact, validates RMFT’s architecture-agnostic design.
Figure 8 compares the average perplexity per checkpoint for both models across all three training techniques.
(a) Average Perplexity for GPT-2-XL
(b) Average Perplexity for GPT-Neo-1.3B
Figure 8: Average perplexity per checkpoint for GPT-2-XL and GPT-Neo-1.3B across all training techniques. (a) GPT-2-XL shows RMFT maintains lower perplexity than deduplication throughout training. (b) GPT-Neo-1.3B exhibits similar trends, with RMFT achieving consistent performance across checkpoints. Both models demonstrate that RMFT preserves model quality better than deduplication.
C.1.3 Summary
These results demonstrate that RMFT’s privacy-utility tradeoff generalizes across model scales and architectures. Both GPT-2-XL (1.5B) and GPT-Neo (1.3B) exhibit:
  • Consistent memorization reduction: approx. 80% average TER and SER reduction
  • Minimal performance impact: less than 8% perplexity increase
  • Similar final checkpoint performance: greater than 82% extraction reduction
This consistency supports RMFT’s applicability to other transformer-based language models.